OpenStudy (masechaba):

Find the plane that passes through the point (2,4,-3) and is parallel to the plane -2x + 4y - 5z + 6 = 0

OpenStudy (anonymous):

let A,B,C are directional ratio of requird plane that passes through (2,4,-3) so eq of plane A(x-xo)+B(y-yo)+C(z-zo)=0 here xo,yo,zo=(2,4,-3) A(x-2)+B(y-4)+C(z+3)=0----->eq1 consider eq of given plane -2x+4y-5z+6=0 A1,B1,C1=-2,4,-5(directional ratios of its normal) since both plane are perallel,thier normal are parallel by parallel line condition A/A1=B/B1=C/C1=K(SAY)

OpenStudy (anonymous):

A=-2K,B=4K,C=-5K PUT IN 1 AND GET THE EQ THATS EQN OF PLANE